Diego Company manufactures one product that is sold for $78 per unit in two geographic regions—the East and West regions. The following information pertains to the company’s first year of operations in which it produced 49,000 units and sold 44,000 units.

Variable costs per unit:
Manufacturing:
Direct materials $ 28
Direct labor $ 14
Variable manufacturing overhead $ 4
Variable selling and administrative $ 6
Fixed costs per year:
Fixed manufacturing overhead $ 686,000
Fixed selling and administrative expense $ 510,000

The company sold 32,000 units in the East region and 12,000 units in the West region. It determined that $230,000 of its fixed selling and administrative expense is traceable to the West region, $180,000 is traceable to the East region, and the remaining $100,000 is a common fixed expense. The company will continue to incur the total amount of its fixed manufacturing overhead costs as long as it continues to produce any amount of its only product.

a. What is the company’s break-even point in unit sales?

b. Is it above or below the actual unit sales?
Answer #1
Variable costs per unit 52 =28+14+4+6
Unit Contribution margin 26 =78-52
Total fixed costs 1196000 =686000+510000
a
Total fixed costs 1196000
Divide by Unit Contribution margin 26
Break-even point in unit sales 46000 units
b
It is above the actual unit sales